# TyphoFormer: Language-Augmented Transformer for Accurate Typhoon Track Forecasting

### v1

@@ Abstract
Accurate typhoon track forecasting is crucial for early system warning and disaster response. While Transformer-based models have demonstrated strong performance in modeling the temporal dynamics of dense trajectories of humans and vehicles in smart cities, they usually lack access to broader contextual knowledge that enhances the forecasting reliability of sparse meteorological trajectories, such as typhoon tracks. To address this challenge, we propose TyphoFormer, a novel framework that incorporates natural language descriptions as auxiliary prompts to improve typhoon trajectory forecasting. For each time step, we use Large Language Model (LLM) to generate concise textual descriptions based on the numerical attributes recorded in the North Atlantic hurricane database. The language descriptions capture high-level meteorological semantics and are embedded as auxiliary special tokens prepended to the numerical time series input. By integrating both textual and sequential information within a unified Transformer encoder, TyphoFormer enables the model to leverage contextual cues that are otherwise inaccessible through numerical features alone. Extensive experiments are conducted on HURDAT2 benchmark, results show that TyphoFormer consistently outperforms other state-of-the-art baseline methods, particularly under challenging scenarios involving nonlinear path shifts and limited historical observations.
1 Introduction
@@
Typhoons and hurricanes pose significant threats to coastal regions, causing devastating damage to infrastructure, ecosystems, and human lives Ramachandran et al. (2025); Wachnicka et al. (2020); Hao et al. (2024). The southeastern region of United States, especially Florida, is facing increasingly frequent and intense storms Oliver-Smith (2019); Malmstadt et al. (2009). For instance, Hurricane IAN (AL092022) in 2022 resulted in over 150 fatalities and an estimated \$113 billion in economic losses, making it one of the costliest hurricanes in U.S. history Wang et al. (2024b); Bushong and Welch (2023). With the rise of sea surface temperatures and the evolution of climate patterns, accurate typhoon trajectory forecasting has become more critical than ever for emergency planning, resource allocation, and early evacuation Wang et al. (2024a). Reliable predictions of storm movement not only support disaster mitigation efforts but also enable authorities to issue timely warnings, minimizing potential casualties and disruption Yu et al. (2022). In light of the substantial socio-economic impact of such extreme weather events, there is a growing demand for advanced, data-driven forecasting methods that can complement or enhance traditional meteorological models.
@@
Existing approaches for typhoon track forecasting generally fall into two main categories: (I) time series modeling methods and (II) physics-based numerical models. To address the need for data-driven forecasting, a growing body of work has explored modeling typhoon tracks as temporal sequences using deep learning techniques Wang et al. (2024c). These methods treat the storm’s historical positions and attributes as a time series and aim to predict its future trajectory based on learned temporal patterns using deep learning-based methods. Early approaches employed recurrent neural network (RNN)-based architectures such as LSTM, GRU, and their variants to capture sequential dependencies in storm movement Gao et al. (2018); Song et al. (2022); Wei and Cheng (2020). More recent advances have leveraged Transformer-based architectures, which offer improved long-range modeling capabilities and greater scalability for sequence forecasting Jiang et al. (2023); Gan et al. (2024). Methods such as InFormer Zhou et al. (2021) and AutoFormer Wu et al. (2021) have demonstrated promising results in related time series domains, particularly for long-horizon prediction tasks. While these models excel in learning from past trajectory data, they lack an awareness of external contextual information, such as meteorological semantics or high-level event dynamics, that may further enhance predictive accuracy.
@@
As a basic tool in traditional meteorological analysis, physics-based numerical models are also widely used in operational typhoon forecasting Iwasaki et al. (1987); Knaff et al. (2007). Numerical models simulate atmospheric dynamics by solving complex partial differential equations derived from physical laws Bauer et al. (2015), such as the Navier–Stokes Equations Rozanova et al. (2010) and Boundary Layer Structure Modeling Kepert (2010). Prominent forecasting systems, including the the Global Forecast System (GFS) Halperin et al. (2020) and Climatology and Persistence Model Knaff et al. (2007), use large-scale initial conditions, such as pressure fields, wind vectors, and sea surface temperatures, to estimate future storm trajectories. While these models offer physical interpretability and have been refined through decades of meteorological research, they require gigantic computational resources (e.g., supercomputing center) and are sensitive to initial condition uncertainties. Furthermore, their performance may degrade in data-sparse regions or during rapidly evolving storm phases, where high-resolution input data are limited or delayed. As a result, complementary deep learning and data-driven techniques have been increasingly explored to address these limitations inherent in traditional approaches.
@@
2 Methodology
We encode both structured meteorological features and textual descriptions into a unified Transformer-based architecture, enabling the model to capture both temporal dynamics and contextual knowledge. The overall model architecture is shown in Figure 1.
2.1 Trajectory Data and Problem Formulation
@@
Figure 1: The overall model architecture of TyphoFormer.
2.2 Language Contextual Prompt Generation
@@
2.3 Text Embedding and Feature Fusion
To integrate semantically rich natural language prompts into the forecasting model, we first encode each textual description into a dense sequence of embeddings using a pretrained language encoder. Specifically, we employ BERT to convert each prompt into a sequence $\mathcal{P}=\{\mathbf{p}_{1},\mathbf{p}_{2},\dots,\mathbf{p}_{M}\}$, where each token embedding $\mathbf{p}_{m}\in\mathbb{R}^{d}$. Meanwhile, the structured time series input is denoted as $\mathcal{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}\}$, with each $\mathbf{x}_{t}\in\mathbb{R}^{d}$ representing multivariate meteorological features at $t$.
@@ 2.3 Text Embedding and Feature Fusion
$\tilde{\mathbf{x}}_{t}=\sigma(W_{g}[\mathbf{x}_{t};\mathbf{p}_{\text{mean}}]+b_{g})\odot\mathbf{x}_{t}+\left(1-\sigma(\cdot)\right)\odot\mathbf{p}_{\text{mean}}$ (1)
@@
2.4 Forecasting Head and Training
@@ 3 Experiments
Baseline methods. We compare our TyphoFormer [^2] https://github.com/LabRAI/TyphoFormer against two categories of baselines. The first category is the classical numerical equation-based models, we choose the representative (I) CLIPER (Climatology and Persistence Model) Knaff et al. (2007), which extrapolates storm trajectories using historical patterns and persistence assumptions. The second category is time-series forecasting models, including recurrent models like (III) LSTM and (IV) GRU, as well as Transformer-based or MLP-based advanced architectures like (V) Informer Zhou et al. (2021), (VI) Autoformer Wu et al. (2021), and (VII) TSMixer Ekambaram et al. (2023). These baselines provide a diverse comparison across classical statistical models and advanced deep learning frameworks.
3.1 Main Results
Table 1: Performance Comparison of Various Methods.
Best values are in bold font, the second-best values are underlined. Models MAE (All) $\Delta$R (km) (All) 6h 12h 18h 24h 6h 12h 18h 24h CLIPER 0.235 0.275 0.310 0.368 34.265 42.205 51.632 58.268 GRU 0.367 0.405 0.493 0.640 50.480 69.397 90.875 103.894 LSTM 0.392 0.431 0.583 0.828 46.096 71.365 95.412 112.663 Informer 0.289 0.318 0.392 0.483 37.592 46.435 56.433 76.684 Autoformer 0.263 0.286 0.357 0.433 39.836 47.183 63.775 70.862 TSMixer 0.214 0.268 0.297 0.353 35.720 45.265 50.330 62.910 TyphoFormer 0.188 0.242 0.261 0.312 31.539 38.084 42.435 49.562 Models MAE (2024) $\Delta$R (km) (2024) 6h 12h 18h 24h 6h 12h 18h 24h CLIPER 0.221 0.269 0.294 0.361 33.915 44.542 48.580 57.642 GRU 0.359 0.394 0.491 0.632 49.674 70.357 92.542 105.114 LSTM 0.388 0.422 0.587 0.835 44.203 69.051 94.555 109.108 Informer 0.290 0.316 0.395 0.487 37.843 48.791 57.445 78.497 Autoformer 0.257 0.284 0.358 0.435 39.128 46.085 65.039 72.827 TSMixer 0.228 0.265 0.305 0.347 35.148 43.935 52.501 63.118 TyphoFormer 0.185 0.241 0.263 0.317 31.274 37.934 42.973 50.881
Figure 2: Trajectory forecasting performance among all the methods on hurricane MILTON 2024.

### v2

@@ Abstract.
Accurate typhoon track forecasting is crucial for early system warning and disaster response. While Transformer-based models have demonstrated strong performance in modeling the temporal dynamics of dense trajectories of humans and vehicles in smart cities, they usually lack access to broader contextual knowledge that enhances the forecasting reliability of sparse meteorological trajectories, such as typhoon tracks. To address this challenge, we propose TyphoFormer, a novel framework that incorporates natural language descriptions as auxiliary prompts to improve typhoon trajectory forecasting. For each time step, we use Large Language Model (LLM) to generate concise textual descriptions based on the numerical attributes recorded in the North Atlantic hurricane database. The language descriptions capture high-level meteorological semantics and are embedded as auxiliary special tokens prepended to the numerical time series input. By integrating both textual and sequential information within a unified Transformer encoder, TyphoFormer enables the model to leverage contextual cues that are otherwise inaccessible through numerical features alone. Extensive experiments are conducted on HURDAT2 benchmark, results show that our TyphoFormer consistently outperforms other state-of-the-art baseline methods, particularly under challenging scenarios involving nonlinear path shifts and limited historical observations.
Typhoon Track Prediction, Geospatial Time Series, Natural Language Prompt, Multimodal Fusion, Large Language Models
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
@@
Typhoons and hurricanes pose significant threats to coastal regions, causing devastating damage to infrastructure, ecosystems, and human lives (Ramachandran et al., 2025; Wachnicka et al., 2020; Hao et al., 2024). The southeastern United States, especially Florida, is facing increasingly frequent and intense storms (Oliver-Smith, 2019; Malmstadt et al., 2009). For instance, Hurricane IAN in 2022 resulted in over 150 fatalities and an estimated \$113 billion in economic losses, making it one of the costliest hurricanes in U.S. history (Wang et al., 2024c; Bushong and Welch, 2023). With the rise of sea surface temperatures and the evolution of climate patterns, accurate typhoon trajectory forecasting has become more critical than ever for emergency planning, resource allocation, and early evacuation (Wang et al., 2024a). Reliable predictions of storm movement not only support disaster mitigation efforts but also enable authorities to issue timely warnings, minimizing potential casualties and disruption (Yu et al., 2022). In light of the substantial socio-economic impact of such extreme weather events, there is a growing demand for advanced, data-driven forecasting methods that can complement or enhance traditional meteorological models.
@@
Existing approaches for typhoon track forecasting generally fall into two main categories: (I) time series modeling methods and (II) physics-based numerical models. To address the need for data-driven forecasting, a growing body of work has explored modeling typhoon tracks as temporal sequences using deep learning techniques (Wang et al., 2024b). These methods treat the storm’s historical positions and attributes as a time series and aim to predict its future trajectory based on learned temporal patterns using deep learning-based methods. Early approaches employed recurrent neural network (RNN)-based architectures such as LSTM, GRU, and their variants to capture sequential dependencies in storm movement (Gao et al., 2018; Song et al., 2022; Wei and Cheng, 2020). More recent advances have leveraged Transformer-based architectures, which offer improved long-range modeling capabilities and greater scalability for sequence forecasting (Jiang et al., 2023; Gan et al., 2024). Methods such as InFormer (Zhou et al., 2021) and AutoFormer (Wu et al., 2021) have demonstrated promising results in related time series domains, particularly for long-horizon prediction tasks. While these models excel in learning from past trajectory data, they lack an awareness of external contextual information, such as meteorological semantics or high-level event dynamics, that may further enhance predictive accuracy.
@@
As a basic tool in traditional meteorological analysis, physics-based numerical models are also widely used in operational typhoon forecasting (Iwasaki et al., 1987; Knaff et al., 2007). Numerical models simulate atmospheric dynamics by solving complex partial differential equations derived from physical laws (Bauer et al., 2015), such as the Navier–Stokes Equations (Rozanova et al., 2010) and Boundary Layer Structure Modeling (Kepert, 2010). Prominent forecasting systems, including the the Global Forecast System (GFS) (Halperin et al., 2020) and Climatology and Persistence Model (Knaff et al., 2007), use large-scale initial conditions, such as pressure fields, wind vectors, and sea surface temperatures, to estimate future storm trajectories. While these models offer physical interpretability and have been refined through decades of meteorological research, they require gigantic computational resources (e.g., supercomputing center) and are sensitive to initial condition uncertainties. Furthermore, their performance may degrade in data-sparse regions or during rapidly evolving storm phases, where high-resolution input data are limited or delayed. As a result, complementary deep learning and data-driven techniques have been increasingly explored to address these limitations inherent in traditional approaches.
@@
2. Methodology
We encode both structured meteorological features and textual descriptions into a unified Transformer-based architecture, enabling the model to capture both temporal dynamics and contextual knowledge. The overall model architecture is illustrated in Figure 1.
2.1. Trajectory Data and Problem Formulation
@@
Figure 1. The overall model architecture of TyphoFormer.
2.2. Language Contextual Prompt Generation
@@
2.3. Text Embedding and Feature Fusion
To integrate semantically rich natural language prompts into the forecasting model, we first encode each textual description into a dense sequence of embeddings using a pretrained language encoder. Specifically, we employ BERT to convert each prompt into a sequence $\mathcal{P}=\{\mathbf{p}_{1},\mathbf{p}_{2},\dots,\mathbf{p}_{M}\}$, where each token embedding $\mathbf{p}_{m}\in\mathbb{R}^{d}$. Meanwhile, the structured time series input is denoted as $\mathcal{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}\}$, with each $\mathbf{x}_{t}\in\mathbb{R}^{d}$ representing multivariate meteorological features at time step $t$.
@@ 2.3. Text Embedding and Feature Fusion
(1) $\tilde{\mathbf{x}}_{t}=\sigma(W_{g}[\mathbf{x}_{t};\mathbf{p}_{\text{mean}}]+b_{g})\odot\mathbf{x}_{t}+\left(1-\sigma(\cdot)\right)\odot\mathbf{p}_{\text{mean}}$
@@
2.4. Forecasting Head and Training
@@ 3. Experiments
Baseline methods. We compare our TyphoFormer [^2] https://github.com/LabRAI/TyphoFormer against two categories of baselines. The first category is the classical numerical equation-based models, we choose the representative (I) CLIPER (Climatology and Persistence Model) (Knaff et al., 2007), which extrapolates storm trajectories using historical patterns and persistence assumptions. The second category is time-series forecasting models, including recurrent models like (III) LSTM and (IV) GRU, as well as Transformer-based or MLP-based advanced architectures like (V) Informer (Zhou et al., 2021), (VI) Autoformer (Wu et al., 2021), and (VII) TSMixer (Ekambaram et al., 2023). These baselines provide a diverse comparison across classical statistical models and advanced deep learning frameworks.
3.1. Main Results
Table 1. Performance Comparison of Various Methods.
The best values are in bold font, the second-best values are indicated with underlining. Models MAE (All) $\Delta$R (km) (All) 6h 12h 18h 24h 6h 12h 18h 24h CLIPER 0.235 0.275 0.310 0.368 34.265 42.205 51.632 58.268 GRU 0.367 0.405 0.493 0.640 50.480 69.397 90.875 103.894 LSTM 0.392 0.431 0.583 0.828 46.096 71.365 95.412 112.663 Informer 0.289 0.318 0.392 0.483 37.592 46.435 56.433 76.684 Autoformer 0.263 0.286 0.357 0.433 39.836 47.183 63.775 70.862 TSMixer 0.214 0.268 0.297 0.353 35.720 45.265 50.330 62.910 TyphoFormer 0.188 0.242 0.261 0.312 31.539 38.084 42.435 49.562 Models MAE (2024) $\Delta$R (km) (2024) 6h 12h 18h 24h 6h 12h 18h 24h CLIPER 0.221 0.269 0.294 0.361 33.915 44.542 48.580 57.642 GRU 0.359 0.394 0.491 0.632 49.674 70.357 92.542 105.114 LSTM 0.388 0.422 0.587 0.835 44.203 69.051 94.555 109.108 Informer 0.290 0.316 0.395 0.487 37.843 48.791 57.445 78.497 Autoformer 0.257 0.284 0.358 0.435 39.128 46.085 65.039 72.827 TSMixer 0.228 0.265 0.305 0.347 35.148 43.935 52.501 63.118 TyphoFormer 0.185 0.241 0.263 0.317 31.274 37.934 42.973 50.881
Figure 2. Trajectory forecasting performance among all the methods on hurricane MILTON 2024.